OpenStudy (anonymous):

Calculus True or False???? Help?? The graph of f(x) =1/x is concave downward for x<0 and concave upward for x>0 and thus it has a point of inflection at x=0.

OpenStudy (amistre64):

how do we find a point of inflection?

OpenStudy (amistre64):

and also, inflection presupposes conitnuous along hte domain i believe

OpenStudy (anonymous):

I'm pretty sure point of inflection is the answers to the second derivative no?

OpenStudy (amistre64):

yes, 2nd derivative helps define critical values for infleciton points

OpenStudy (amistre64):

but inflection points tend to require a continuous function, at least continuous around the interval that is getting inflected

OpenStudy (anonymous):

So I've been doing this very wrong. :( Anyways, how do I know if this is true or not??

OpenStudy (anonymous):

And after this question could I post a function that needs inflection points found? Because from what you said, I think I did it wrong.

OpenStudy (anonymous):

@mathmale @abb0t could you help me with this?

OpenStudy (anonymous):

@eliassaab

OpenStudy (ranga):

The domain of the function f(x) = 1/x excludes x = 0. The function is not defined for x = 0 and therefore we cannot discuss the extrema or the point of inflection at x = 0.

OpenStudy (ranga):

If a function is continuous over an interval and it changes from concave down to concave up or vice-versa in that interval then it will have a point of inflection in that interval. Here 1/x is concave down for x < 0 and concave up for x > 0 but at x = 0 the function is not defined and therefore x = 0 is not a point of inflection because f(x) is not continuous at x = 0.

OpenStudy (loser66):

This site said that f"(0) >0 for x>0 and f"(0) for x<0 shows that the point 0 is inflection point although f"(0) does not exist. http://www.sosmath.com/calculus/diff/der15/der15.html at the middle of the page

OpenStudy (ranga):

@Loser66 That is a slightly different point. For c to be a possible point of inflection, f''(c) = 0 or f''(c) does not exist. But f(x) should be continuous and differentiable at x = c. In that link they tell you that the function is continuous and differentiable at all x. Here, f(x) = 1/x. f''(0) does not exist. Is x = 0 an inflection point? Since x = 0 is not in the domain of f(x), the function is not defined at x = 0 and so it cannot have a point of inflection at x = 0.

OpenStudy (loser66):

Yes, got you, thanks @ranga

OpenStudy (anonymous):

@ranga thank you so much